# Water Ice in the Edge-On Orion Silhouette Disk 114–426 from JWST NIRCam Images

### v1

@@ Abstract
We examine images of the protoplanetary disk 114–426 with JWST/NIRCam in 12 bands. This large disk is oriented edge-on with a dark midplane flanked by lobes of scattered light. The outer edges of the midplane are seen in silhouette against the Orion Nebula, providing a unique opportunity to study planet-forming material in absorption. We discover a dip in the scattered light of the disk at 3 – compelling evidence for the presence of water ice. The 3 dip is also seen in the silhouette of the disk, where we quantify the ice abundance with models of pure absorption and avoid the complications of disk scattering effects. We find grain ice-to-refractory mass ratios of up to 0.2, maximum grain sizes of 0.25 to 5 , and a total dust plus ice mass of 0.46 in the silhouette region. We also discover excess absorption in the NIRCam bands that include the Paschen line, suggesting there may be excited atomic hydrogen in the disk. Examining the morphology of the scattered light lobes reveals that they are laterally offset from each other and exhibit a brightness asymmetry that flips with wavelength – both evidence for a tilted inner disk in this system.
@@ 1 INTRODUCTION
The 3 feature has been observed towards a number of edge-on protoplanetary disks (Pontoppidan et al., 2005; Terada et al., 2007; McCabe et al., 2011; Sturm et al., 2023a). In these systems, the scattered light from the star and warm inner disk serves as the background continuum on which the ice features are imprinted. The scattering of radiation through the disk complicates the interpretation of these observations (Ballering et al., 2021; Sturm et al., 2023b) such that the column density of ice cannot be directly inferred from the depth of the absorption feature, as is possible in the case of pure absorption against a background source.
@@
The disk 114–426 provides a unique opportunity to study planet-forming material in absorption. 114–426, named using the coordinate-based convention introduced by O’Dell & Wen (1994), resides in the dense Trapezium Cluster, itself part of the wider Orion Nebula Cluster (ONC). It is located to the southwest of the Trapezium OB stars. The disk has a major axis of , corresponding to a diameter of 1,053 au at a distance of 390 pc (Maíz Apellániz et al., 2022). It is oriented edge-on, as evident by the dark lane of the disk midplane that occults the central star, with surrounding lobes of scattered light. Additionally, the outer ansae of the dark lane extend beyond the radial extent of the scattered light and are seen in silhouette against the emission of the background Orion Nebula. This provides a region where the disk can be studied in pure absorption.
@@
114–426 has been the subject of several detailed imaging studies at optical and near-IR wavelengths, primarily with the Hubble Space Telescope (McCaughrean & O’Dell, 1996; McCaughrean et al., 1998; Bally et al., 2000; Throop et al., 2001; Shuping et al., 2003; Miotello et al., 2012). It was realized that the outer ansae of the disk are somewhat transparent to the background nebula at optical wavelengths. Examining the variation in optical depth with wavelength provides a means to measure the dust column density and grain sizes. The most recent analysis used absorption measurements in five visible wavelength bands and constrained the radius of the grains to be 0.2–0.7 (Miotello et al., 2012).
@@
ALMA observations of 114–426 detected the sub-mm continuum emission from large dust grains in the disk, while the sub-mm CO lines were seen in absorption against the nebular background (Mann et al., 2014; Bally et al., 2015). The velocity profile of the CO measurement constrained the stellar mass to 0.4–1 , making this disk’s large outer extent especially remarkable.
@@
In the present paper, we analyze James Webb Space Telescope (JWST) NIRCam images of the 114–426 disk in 12 wide-, medium-, and narrow-band filters spanning 1–5 . These images show convincing evidence for the 3 water ice absorption feature in both (1) lines-of-sight dominated by scattered light, as well as (2) the outer ansae of the disk detected in silhouette against the bright background nebula. This analysis provides the first demonstration of NIRCam’s ability to shed light on the nature of water ice in young circumstellar disks and provides some of the strongest direct constraints on the amount of water ice at large disk radii given the disk’s unique orientation and position in the ONC.
@@ 2 Methods
JWST/NIRCam imaged a large region () of the ONC, including 114–426, for Cycle 1 GTO Program 1256 (PI McCaughrean). For a complete overview of these observations see McCaughrean & Pearson (2023). Twelve filters were used spanning the 1–5 wavelength range: F115W, F140M, F162M, F182M, F187N, F212N, F277W, F300M, F335M, F360M, F444W, and F470N. The first six of these are from the Short Wavelength (SW) channel and the other six from the Long Wavelength (LW) channel.
We use the customized pipeline reduction of the images described by McCaughrean & Pearson (2023). While the Level 2 pipeline-processed individual images available directly from the Mikulski Archive for Space Telescopes (MAST) database yielded reasonable results, the default Level 3 processing to combine the many images into large-scale mosaics did not. In addition to astrometric errors, the default pipeline did not result in good matching of the bright and variable nebulosity across the Orion Nebula, leading to large-scale intensity steps and thus erroneous nebular fluxes. The custom pipeline processing corrected for the astrometric offsets and delivered smoothly-matched large-scale mosaics.
@@
One factor that has complicated prior analyses of the 114–426 disk is that the telescope PSF blurs bright emission from the surrounding nebula and the scattered light lobes into the silhouette, potentially leading to an overestimation of the transmitted flux (McCaughrean & O’Dell, 1996). Because the diffraction-limited PSF size increases with wavelength, this could bias the observed slope of the absorption spectra.
@@
We estimate the pixel-to-pixel noise in each image by computing the standard deviation of brightness values in a rectangular aperture offset from the disk. For the unconvolved images, these noise levels are: 2.5, 3.0, 3.2, 2.8, 2.8, 3.7, 3.6, 4.5, 10.1, 6.8, 4.7, and 4.8 percent of the median flux in the region in the 12 bands in order of increasing wavelength. When comparing the brightness between bands, we also include a 2% flux calibration uncertainty. This is conservative, as most bands have 2% uncertainty, according to the current online documentation.
@@ 3.1 Disk Morphology
Figure 1 shows detailed images of 114–426. The left panel is a color composite of the four wide- and medium-width SW bands plus F277W, and the right panel shows the F187N band. In both images, pixel scale is oversampled and the intensities are scaled logarithmically to compress the dynamic range and enhance fainter details. Figure 2 displays a gallery of the 12 NIRCam images of 114–426, this time all sampled to the LW pixel scale for more direct comparison and displayed with a linear brightness scale. Figure 3 presents surface brightness cuts along the major and minor axes of the disk.
The images reveal a bi-lobed emission pattern surrounding a dark lane and flanked by even darker extensions seen in silhouette against the nebular background. The dark ansae curve away from the midplane in opposite directions, giving the disk a warped S-shaped structure. The ends of the ansae appear wispy with the northeast silhouette being more extended and having a hook-like shape. To first-order, this morphology is consistent with that seen in HST images (McCaughrean et al., 1998; Miotello et al., 2012).
The F187N image differs from the other bands. This narrow-band filter is centred on the Pa- emission line at 1.875, thus admitting significant light from the background H II and PDR while minimizing any continuum contribution. The scattered light lobes are barely visible, and the whole radial and vertical extent of the disk appears in silhouette. A similar appearance is evident in HST optical images that trace the H line (McCaughrean et al., 1998), not least because the diffraction-limit of the images is very similar (the 2.7 times larger diameter of the JWST primary mirror offsetting the 2.9 times longer wavelength of the Pa- line.)
A detailed examination of the scattered light lobes shows a wavelength-dependent brightness asymmetry between them. At wavelengths shorter than 2 , the northwest lobe is brighter, while at wavelengths longer than 2 the asymmetry flips and the southeast lobe is brighter. This behavior was seen with HST/NICMOS when comparing images at 0.57, 1.1, 1.6, and 2.0 (McCaughrean et al., 1998), and here we find that the trend continues out to 4.7 . The pattern is clear from the redder color of the southeast lobe in the left panel of Figure 1 and from the minor axis cuts shown in Figure 3. A similar behavior is seen in edge-on disks around 2MASS J16281370-2431391 (a.k.a. The Flying Saucer; Grosso et al., 2003) and IRAS 04302+2247 (a.k.a. The Butterfly Star; Villenave et al., 2024), although in the latter case the brightness flip occurs at longer wavelengths between 12.8 and 21.1 .
@@ 3.2.1 Silhouette Region
At the shortest three wavelengths (F115W, F140M, and F162M) the disk becomes increasingly transparent with increasing wavelengths. This is the expected behavior of small ( few ) dust grains and continues the trend seen from shorter wavelengths with HST (Miotello et al., 2012). We quantify the grain sizes in Section 3.3. At the darkest parts of the silhouette, 20–40% of the background light is transmitted at these wavelengths.
The next two bands (F182M and F187N) show a dip in the absorption spectra at many of the pixels. These bands include the Pa- line, in which the background is very bright. The fact that the disk absorbs more photons at this wavelength than at adjacent wavelengths is unexpected and cannot be explained by the opacity of bare or icy dust grains. This suggests that the gas in the disk contributes to the opacity in these bands. We discuss this possibility more in Section 4.2.
Another prominent dip in the absorption spectra is seen in the F300M band. This filter has a medium bandwidth and does not sample a bright line from the nebular background. Thus, we associate the feature with water ice in the disk, as seen in many other protostellar and protoplanetary systems (see Section 1.) The 3 water ice feature is relatively broad, but its shape is not fully sampled by the available photometry. The F300M band does not probe the full depth of the feature, and the F277W and F335M bands likely sample of mix of the wings of the feature and the neighboring continuum.
At the longest wavelengths (F360M, F444W, F470N) the absorption spectra continue to rise, with the disk becoming nearly totally transparent at the outer edge of the ansae (e.g., locations A and D).
@@ 3.2.2 Scattered Light Region
In Figure 5 we show the spectra at various locations centered on the scattered light lobes (U, V, Y, and Z) and in the dark lane between them (W and X). Because the disk here is not seen in absorption relative to the background, we do not divide the brightness by the background. The error bars include the measured statistical uncertainty plus the 2% calibration uncertainty. As with Figure 4, the black spectra are extracted from the unconvolved images while gray spectra are extracted from the convolved images, and the difference is discussed in the Appendix.
In these locations, the Pa- line is evident in emission from an excess in the F187N band. The amplitude does not trace the true intensity of the line because it is diluted by the width of the band, which is likely wider than that of the line. Interpreting the source of these Pa- photons is difficult. They could arise from the background nebula and are partially transmitted through this region of the disk, although unlike in the silhouette, the optical depth of the scattered light lobes cannot easily be quantified. Other potential sources of these Pa- photons are accretion onto the central star with the photons then scattering out of the disk or emission from excited hydrogen gas in the disk. Future spectroscopic studies covering additional nebular lines could test the degree to which the lines in the scattered light lobes resemble those from the background. Spectra of the scattered light lobes would also be useful to constrain the spectral type of the central star.
A dip in the spectra from water ice at 3 is evident in this region of the disk as well. This is consistent with predictions from radiative transfer models of ice-bearing edge-on disks (Ballering et al., 2021; Sturm et al., 2023b) and from observations of such disks (e.g., Sturm et al., 2023a). The brightness of the disk in the F300M filter – even in the center of the dark lane – does not go to zero. This is consistent with measurements of the 3 feature in the dark lane of other disks such as HH48 NE (Sturm et al., 2023a). One possible explanation is that photons from ice-free regions of the disk scatter into the same lines of sight as those from icy regions. This adds a floor of continuum at 3 , preventing the flux from dropping to zero even if the disk midplane is actually very ice-rich. It could also be due to the NIRCam PSF blurring signal into the midplane from the bright lobes. Furthermore, PAH emission from the disk at 3.3 could influence the shape of the feature, as was observed in the spectrum of the edge-on disk HH48 NE (Sturm et al., 2023a, 2024). Interpreting the details of the spectra in the scattered light region requires fitting the data with radiative transfer models, which is beyond the scope of this study.
@@ 3.3 Quantifying the Water Ice Abundance, Grain Sizes, and Dust Mass
We now return our attention to the silhouette of the disk. Given these lines-of-sight primarily intercept material at the very edges of the disk under more similar physical and chemical conditions compared to lines-of-sight directly toward the central star, a relatively straightforward approach can be used to extract the water ice-to-dust composition as described in the following paragraphs.
@@
where is the dust absorption opacity (cm2 g-1) and is the dust surface density (g cm-2). reflects the amount of dust along the line of sight and is independent of wavelength, while varies with wavelength and is set by the grain size and composition. We choose to include only the absorption opacity and not the scattering opacity because scattering will be preferentially in the forward direction and thus photons from the background will mostly continue into our line-of-sight. The model assumes there is no emission from the silhouette of the disk or from nebular (or interstellar) material between the disk and Earth. This is supported by the conclusion that silhouette disks like 114–426 likely reside in the foreground of the Orion Nebula (McCaughrean & O’Dell, 1996; Miotello et al., 2012).
We generate model opacity spectra using the code optool (Dominik et al., 2021). The minimum grain size is fixed at 0.005 , while the maximum grain size is allowed to vary as a free parameter, and the size distribution is a power-law with an index of as is typically assumed. The grains are modeled as a refractory core with an ice mantle. The core composition is set to a mixture of 87% amorphous pyroxene (with a 70% Mg content) and 13% amorphous carbon by mass. This is a commonly-assumed core composition (Woitke et al., 2016) and a default option in optool. The mass of the water ice mantle is parameterized as a fraction of the core mass. The optical constants for pyroxene, carbon, and water ice are from Dorschner et al. (1995), Zubko et al. (1996), and Warren & Brandt (2008), respectively. Note that optical constants of ice vary by up to 15% depending on the temperature and porosity (e.g., Rocha et al., 2024).
There are thus three free parameters in our model: (1) the ice-to-refractory mass ratio, (2) the maximum grain size, and (3) . The first two parameters enter Equation 1 through . We generate model absorption spectra on a finely-sampled wavelength grid then compute synthetic photometry for the 12 NIRCam bands by integrating the model weighted by the band transmission profiles. We do not include all bands in the fit. We exclude the F182M and F187N bands because there is apparent excess absorption of Pa- photons in these bands, which is not included in our simplified model. We also exclude the other two narrow-band filters in the set: F212N and F470N. These are centered on molecular H2 transitions and thus could potentially be contaminated by absorbing gas in the disk. We use a grid search to find the model that best fits the remaining eight bands by minimizing the metric:
where the sum is over bands included in the fit. We fit each pixel in the silhouette of the disk independently. We fit pixels where the absorption depth in the shortest wavelength band is 0.9 and where the scattered light does not contaminate the absorption spectrum. We also fit these models to the spatially-convolved images, and the results are discussed in the Appendix.
@@
By summing the surface density from each pixel, we can derive the total mass of ice and dust in the silhouettes. At a distance of 390 pc, each pixel corresponds to a physical area of cm2. We find 0.27 of dust and ice in the northeast ansa and 0.19 in the southwest ansa, totaling 0.46 of solid material in the silhouette. Bally et al. (2015) find a dust mass of 9.9 in the scattered-light region of the disk from ALMA continuum measurements. So while the silhouette occupies a sizeable fraction of the disk’s area on the sky, it harbors a smaller fraction of the total mass because it lacks larger mm-size grains where most of the solid mass is sequestered.
We show the spectral fit at three pixels in Figure 7. These pixels are at the locations labeled A, B, and C in Figure 4. The values of the best-fit model parameters, along with their uncertainties, are noted in the figure. These locations illustrate a range of ice mass ratios seen across the northeast ansa. The quality of these fits is typical for the pixels throughout the silhouette. The model fits the rising absorption spectrum in the shortest three bands well. The model greatly over-predicts the F182M and F187N bands (excluded from the fit) that may be impacted by preferential absorption of Pa- photons. The model also sometimes over-predicts the data in the excluded F212N band. The F277W and F300M points are fit well. F300M samples the core of the ice absorption feature (although the photometry does not trace the full depth of the feature seen in the high-resolution model), while F277W and F335M are affected by the wings of the feature. F335M is regularly over-predicted by the model. The final two bands that we fit, F360M and F444W, show varying degrees of agreement with the model. Overall this suggests that the shape of the red edge of the water feature is not reproduced correctly in our model, or that there are other sources of opacity that we have not considered. We discuss this further in Section 4.1.
Figure 8 demonstrates the impact of each of the three free parameters on the model behavior. The depth of the 3 feature is highly sensitive to even small water ice fractions (left panel). High ice fractions also shift the model vertically at wavelengths outside of the feature, but such shifts are likely degenerate with the other model parameters. Increasing the maximum grain size flattens the slope of the model in the 1–3 region (middle panel). The variation is most drastic between sizes 0.2 and 5 with very little variation among the models at smaller sizes. At sizes larger than 5 the slope remains constant but the models shift vertically. This is degenerate with other model parameters, which is why we limit the fits to maximum grain sizes 5 . Varying the solid surface density scales the entire absorption spectrum vertically, as expected (right panel).

### v2

@@ Abstract
We examine images of the protoplanetary disk 114–426 with JWST/NIRCam in 12 bands. This large disk is oriented edge on with a dark midplane flanked by lobes of scattered light. The outer edges of the midplane are seen in silhouette against the Orion Nebula, providing a unique opportunity to study planet-forming material in absorption. We discover a dip in the scattered light of the disk at 3 – compelling evidence for the presence of water ice. The 3 dip is also seen in the silhouette of the disk, where we quantify the ice abundance with models of pure absorption and avoid the complications of disk scattering effects. We find grain ice-to-refractory mass ratios of up to 0.2, maximum grain sizes of 0.25–5 , and a total dust plus ice mass of 0.46 in the silhouette region. We also discover excess absorption in the NIRCam bands that include the Pa line, suggesting there may be excited atomic hydrogen in the disk. Examining the morphology of the scattered light lobes reveals that they are laterally offset from each other and exhibit a brightness asymmetry that flips with wavelength – both evidence for a tilted inner disk in this system.
@@ 1 INTRODUCTION
The 3 feature has been observed toward a number of edge-on protoplanetary disks (Pontoppidan et al., 2005; Terada et al., 2007; McCabe et al., 2011; Sturm et al., 2023a). In these systems, the scattered light from the star and warm inner disk serves as the background continuum on which the ice features are imprinted. The scattering of radiation through the disk complicates the interpretation of these observations (Ballering et al., 2021; Sturm et al., 2023b) such that the column density of ice cannot be directly inferred from the depth of the absorption feature, as is possible in the case of pure absorption against a background source.
@@
The disk 114–426 provides a unique opportunity to study planet-forming material in absorption. 114–426, named using the coordinate-based convention introduced by O’Dell & Wen (1994), resides in the dense Trapezium Cluster, itself part of the wider Orion Nebula Cluster (ONC). It is located to the southwest of the Trapezium OB stars. The disk has a major axis of , corresponding to a diameter of 1053 au at a distance of 390 pc (Maíz Apellániz et al., 2022). It is oriented edge on, as evident by the dark lane of the disk midplane that occults the central star, with surrounding lobes of scattered light. Additionally, the outer ansae of the dark lane extend beyond the radial extent of the scattered light and are seen in silhouette against the emission of the background Orion Nebula. This provides a region where the disk can be studied in pure absorption.
@@
114–426 has been the subject of several detailed imaging studies at optical and near-IR wavelengths, primarily with the Hubble Space Telescope (HST; McCaughrean & O’Dell, 1996; McCaughrean et al., 1998; Bally et al., 2000; Throop et al., 2001; Shuping et al., 2003; Miotello et al., 2012). It was realized that the outer ansae of the disk are somewhat transparent to the background nebula at optical wavelengths. Examining the variation in optical depth with wavelength provides a means to measure the dust column density and grain sizes. The most recent analysis used absorption measurements in five visible wavelength bands and constrained the radius of the grains to be 0.2–0.7 (Miotello et al., 2012).
@@
Atacama Large Millimeter/submillimeter Array (ALMA) observations of 114–426 detected the submillimeter continuum emission from large dust grains in the disk, while the submillimeter CO lines were seen in absorption against the nebular background (Mann et al., 2014; Bally et al., 2015). The velocity profile of the CO measurement constrained the stellar mass to 0.4–1 , making this disk’s large outer extent especially remarkable.
@@
In the present paper, we analyze James Webb Space Telescope (JWST) NIRCam images of the 114–426 disk in 12 wide-, medium-, and narrowband filters spanning 1–5 . These images show convincing evidence for the 3 water ice absorption feature in both (1) lines of sight dominated by scattered light as well as (2) the outer ansae of the disk detected in silhouette against the bright background nebula. This analysis provides the first demonstration of NIRCam’s ability to shed light on the nature of water ice in young circumstellar disks and provides some of the strongest direct constraints on the amount of water ice at large disk radii given the disk’s unique orientation and position in the ONC.
@@ 2 Methods
JWST/NIRCam imaged a large region () of the ONC, including 114–426, for Cycle 1 GTO Program 1256 (PI: McCaughrean). For a complete overview of these observations see McCaughrean & Pearson (2023). Twelve filters were used spanning the 1–5 wavelength range: F115W, F140M, F162M, F182M, F187N, F212N, F277W, F300M, F335M, F360M, F444W, and F470N. The first six of these are from the short wavelength (SW) channel and the other six from the long wavelength (LW) channel.
We use the customized pipeline reduction of the images described by McCaughrean & Pearson (2023). While the Level 2 pipeline-processed individual images available directly from the Mikulski Archive for Space Telescopes database yielded reasonable results, the default Level 3 processing to combine the many images into large-scale mosaics did not. In addition to astrometric errors, the default pipeline did not result in good matching of the bright and variable nebulosity across the Orion Nebula, leading to large-scale intensity steps and thus erroneous nebular fluxes. The custom pipeline processing corrected for the astrometric offsets and delivered smoothly matched large-scale mosaics.
@@
One factor that has complicated prior analyses of the 114–426 disk is that the telescope point-spread function (PSF) blurs bright emission from the surrounding nebula and the scattered light lobes into the silhouette, potentially leading to an overestimation of the transmitted flux (McCaughrean & O’Dell, 1996). Because the diffraction-limited PSF size increases with wavelength, this could bias the observed slope of the absorption spectra.
@@
We estimate the pixel-to-pixel noise in each image by computing the standard deviation of brightness values in a rectangular aperture offset from the disk. For the unconvolved images, these noise levels are 2.5%, 3.0%, 3.2%, 2.8%, 2.8%, 3.7%, 3.6%, 4.5%, 10.1%, 6.8%, 4.7%, and 4.8% of the median flux in the region in the 12 bands in order of increasing wavelength. When comparing the brightness between bands, we also include a 2% flux calibration uncertainty. This is conservative, as most bands have 2% uncertainty, according to the current online documentation.
@@ 3.1 Disk Morphology
Figure 1 shows detailed images of 114–426. The left panel is a color composite of the four wide- and medium-width SW bands plus F277W, and the right panel shows the F187N band. In both images, the pixel scale is oversampled and the intensities are scaled logarithmically to compress the dynamic range and enhance fainter details. Figure 2 displays a gallery of the 12 NIRCam images of 114–426, this time all sampled to the LW pixel scale for a more direct comparison and displayed with a linear brightness scale. Figure 3 presents surface brightness cuts along the major and minor axes of the disk.
The images reveal a bilobed emission pattern surrounding a dark lane and flanked by even darker extensions seen in silhouette against the nebular background. The dark ansae curve away from the midplane in opposite directions, giving the disk a warped S-shaped structure. The ends of the ansae appear wispy with the northeast silhouette being more extended and having a hook-like shape. To first order, this morphology is consistent with that seen in HST images (McCaughrean et al., 1998; Miotello et al., 2012).
The F187N image differs from the other bands. This narrowband filter is centered on the Pa emission line at 1.875, thus admitting significant light from the background H II and photodissociation region (PDR) while minimizing any continuum contribution. The scattered light lobes are barely visible, and the whole radial and vertical extent of the disk appears in silhouette. A similar appearance is evident in HST optical images that trace the H line (McCaughrean et al., 1998), not least because the diffraction limit of the images is very similar (the 2.7 larger diameter of the JWST primary mirror offsetting the 2.9 longer wavelength of the Pa line.)
A detailed examination of the scattered light lobes shows a wavelength-dependent brightness asymmetry between them. At wavelengths shorter than 2 , the northwest lobe is brighter, while at wavelengths longer than 2 the asymmetry flips and the southeast lobe is brighter. This behavior was seen with HST/NICMOS when comparing images at 0.57, 1.1, 1.6, and 2.0 (McCaughrean et al., 1998), and here we find that the trend continues out to 4.7 . The pattern is clear from the redder color of the southeast lobe in the left panel of Figure 1 and from the minor axis cuts shown in Figure 3. A similar behavior is seen in edge-on disks around 2MASS J16281370-2431391 (aka The Flying Saucer; Grosso et al., 2003) and IRAS 04302+2247 (aka The Butterfly Star; Villenave et al., 2024), although in the latter case the brightness flip occurs at longer wavelengths between 12.8 and 21.1 .
@@ 3.2.1 Silhouette Region
At the shortest three wavelengths (F115W, F140M, and F162M) the disk becomes increasingly transparent with increasing wavelengths. This is the expected behavior of small ( few micrometer) dust grains and continues the trend seen from shorter wavelengths with HST (Miotello et al., 2012). We quantify the grain sizes in Section 3.3. At the darkest parts of the silhouette, 20%–40% of the background light is transmitted at these wavelengths.
The next two bands (F182M and F187N) show a dip in the absorption spectra at many of the pixels. These bands include the Pa line, in which the background is very bright. The fact that the disk absorbs more photons at this wavelength than at adjacent wavelengths is unexpected and cannot be explained by the opacity of bare or icy dust grains. This suggests that the gas in the disk contributes to the opacity in these bands. We discuss this possibility more in Section 4.2.
Another prominent dip in the absorption spectra is seen in the F300M band. This filter has a medium bandwidth and does not sample a bright line from the nebular background. Thus, we associate the feature with water ice in the disk, as seen in many other protostellar and protoplanetary systems (see Section 1.) The 3 water ice feature is relatively broad, but its shape is not fully sampled by the available photometry. The F300M band does not probe the full depth of the feature, and the F277W and F335M bands likely sample a mix of the wings of the feature and the neighboring continuum.
At the longest wavelengths (F360M, F444W, and F470N) the absorption spectra continue to rise, with the disk becoming nearly totally transparent at the outer edge of the ansae (e.g., locations A and D).
@@ 3.2.2 Scattered Light Region
In Figure 5 we show the spectra at various locations centered on the scattered light lobes (U, V, Y, and Z) and in the dark lane between them (W and X). Because the disk here is not seen in absorption relative to the background, we do not divide the brightness by the background. The error bars include the measured statistical uncertainty plus the 2% calibration uncertainty. As with Figure 4, the black spectra are extracted from the unconvolved images while the gray spectra are extracted from the convolved images, and the difference is discussed in the Appendix.
In these locations, the Pa line is evident in emission from an excess in the F187N band. The amplitude does not trace the true intensity of the line because it is diluted by the width of the band, which is likely wider than that of the line. Interpreting the source of these Pa photons is difficult. They could arise from the background nebula and are partially transmitted through this region of the disk, although unlike in the silhouette, the optical depth of the scattered light lobes cannot easily be quantified. Other potential sources of these Pa photons are accretion onto the central star with the photons then scattering out of the disk or emission from excited hydrogen gas in the disk. Future spectroscopic studies covering additional nebular lines could test the degree to which the lines in the scattered light lobes resemble those from the background. Spectra of the scattered light lobes would also be useful to constrain the spectral type of the central star.
A dip in the spectra from water ice at 3 is evident in this region of the disk as well. This is consistent with predictions from radiative transfer models of ice-bearing edge-on disks (Ballering et al., 2021; Sturm et al., 2023b) and from observations of such disks (e.g., Sturm et al., 2023a). The brightness of the disk in the F300M filter – even in the center of the dark lane – does not go to zero. This is consistent with measurements of the 3 feature in the dark lane of other disks such as HH48 NE (Sturm et al., 2023a). One possible explanation is that photons from ice-free regions of the disk scatter into the same lines of sight as those from icy regions. This adds a floor of continuum at 3 , preventing the flux from dropping to zero even if the disk midplane is actually very ice rich. It could also be due to the NIRCam PSF blurring signal into the midplane from the bright lobes. Furthermore, PAH emission from the disk at 3.3 could influence the shape of the feature, as was observed in the spectrum of the edge-on disk HH48 NE (Sturm et al., 2023a, 2024). Interpreting the details of the spectra in the scattered light region requires fitting the data with radiative transfer models, which is beyond the scope of this study.
@@ 3.3 Quantifying the Water Ice Abundance, Grain Sizes, and Dust Mass
We now return our attention to the silhouette of the disk. Given these lines of sight primarily intercept material at the very edges of the disk under more similar physical and chemical conditions compared to lines of sight directly toward the central star, a relatively straightforward approach can be used to extract the water-ice-to-dust composition as described in the following paragraphs.
@@
where is the dust absorption opacity (units of cm2 g-1) and is the dust surface density (units of g cm-2). reflects the amount of dust along the line of sight and is independent of wavelength, while varies with wavelength and is set by the grain size and composition. We choose to include only the absorption opacity and not the scattering opacity because scattering will be preferentially in the forward direction and thus photons from the background will mostly continue into our line of sight. The model assumes there is no emission from the silhouette of the disk or from nebular (or interstellar) material between the disk and Earth. This is supported by the conclusion that silhouette disks like 114–426 likely reside in the foreground of the Orion Nebula (McCaughrean & O’Dell, 1996; Miotello et al., 2012).
We generate model opacity spectra using the code optool (Dominik et al., 2021). The minimum grain size is fixed at 0.005 , while the maximum grain size is allowed to vary as a free parameter and the size distribution is a power law with an index of as is typically assumed. The grains are modeled as a refractory core with an ice mantle. The core composition is set to a mixture of 87% amorphous pyroxene (with a 70% Mg content) and 13% amorphous carbon by mass. This is a commonly assumed core composition (Woitke et al., 2016) and a default option in optool. The mass of the water ice mantle is parameterized as a fraction of the core mass. The optical constants for pyroxene, carbon, and water ice are from Dorschner et al. (1995), Zubko et al. (1996), and Warren & Brandt (2008), respectively. Note that optical constants of ice vary by up to 15% depending on the temperature and porosity (e.g., Rocha et al., 2024).
There are thus three free parameters in our model: (1) the ice-to-refractory mass ratio, (2) the maximum grain size, and (3) . The first two parameters enter Equation 1 through . We generate model absorption spectra on a finely sampled wavelength grid then compute synthetic photometry for the 12 NIRCam bands by integrating the model weighted by the band transmission profiles. We do not include all bands in the fit. We exclude the F182M and F187N bands because there is apparent excess absorption of Pa photons in these bands, which is not included in our simplified model. We also exclude the other two narrowband filters in the set: F212N and F470N. These are centered on molecular H2 transitions and thus could potentially be contaminated by absorbing gas in the disk. We use a grid search to find the model that best fits the remaining eight bands by minimizing the metric
where the sum is over bands included in the fit. We fit each pixel in the silhouette of the disk independently. We fit pixels where the absorption depth in the shortest wavelength band is 0.9 and where the scattered light does not contaminate the absorption spectrum. We also fit these models to the spatially convolved images, and the results are discussed in the Appendix.
@@
By summing the surface density from each pixel, we can derive the total mass of ice and dust in the silhouettes. At a distance of 390 pc, each pixel corresponds to a physical area of cm2. We find 0.27 of dust and ice in the northeast ansa and 0.19 in the southwest ansa, totaling 0.46 of solid material in the silhouette. Bally et al. (2015) find a dust mass of 9.9 in the scattered light region of the disk from ALMA continuum measurements. So while the silhouette occupies a sizeable fraction of the disk’s area on the sky, it harbors a smaller fraction of the total mass because it lacks larger millimeter-size grains where most of the solid mass is sequestered.
We show a spectral fit at three pixels in Figure 7. These pixels are at the locations labeled A, B, and C in Figure 4. The values of the best-fit model parameters, along with their uncertainties, are noted in the figure. These locations illustrate a range of ice mass ratios seen across the northeast ansa. The quality of these fits is typical for the pixels throughout the silhouette. The model fits the rising absorption spectrum in the shortest three bands well. The model greatly overpredicts the F182M and F187N bands (excluded from the fit) that may be impacted by preferential absorption of Pa photons. The model also sometimes overpredicts the data in the excluded F212N band. The F277W and F300M points are fit well. F300M samples the core of the ice absorption feature (although the photometry does not trace the full depth of the feature seen in the high-resolution model), while F277W and F335M are affected by the wings of the feature. F335M is regularly overpredicted by the model. The final two bands that we fit, F360M and F444W, show varying degrees of agreement with the model. Overall this suggests that the shape of the red edge of the water feature is not reproduced correctly in our model, or that there are other sources of opacity that we have not considered. We discuss this further in Section 4.1.
Figure 8 demonstrates the impact of each of the three free parameters on the model behavior. The depth of the 3 feature is highly sensitive to even small water ice fractions (left panel). High ice fractions also shift the model vertically at wavelengths outside of the feature, but such shifts are likely degenerate with the other model parameters. Increasing the maximum grain size flattens the slope of the model in the 1–3 region (middle panel). The variation is most drastic between sizes of 0.2 and 5 with very little variation among the models at smaller sizes. At sizes larger than 5 the slope remains constant but the models shift vertically. This is degenerate with other model parameters, which is why we limit the fits to maximum grain sizes 5 . Varying the solid surface density scales the entire absorption spectrum vertically, as expected (right panel).